How high (in ft.) will a ball rise if is tossed vertically upward with a speed of 100 ft/s? Hint: 1 inch = 2.54 cm.
1
Expert's answer
2020-11-06T10:13:57-0500
"v^2=2gh\\(100)^2=2(0.0003048)h\\h=1.64\cdot10^7ft"